# Semi-empirical Modeling of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

### v1

@@ Abstract
The evolution of the supermassive Black Hole (BH) population across cosmic times remains a central unresolved issue in modern astrophysics, due to the many noticeable uncertainties in the involved physical processes that span a huge range of spatial, temporal and energy scales. Here we tackle the problem via a semi-empirical approach with minimal assumptions and data-driven inputs. This is firmly grounded on a continuity plus Smoluchowski equation framework that allows to unitarily describe the two primary modes of BH growth: gas accretion and binary mergers. Key quantities related to the latter processes are incorporated through educated parameterizations, and then constrained in a Bayesian setup from joint observational estimates of the local BH mass function, of the large-scale BH clustering, and of the nano-Hz stochastic gravitational wave (GW) background measured from Pulsar Timimg Array (PTA) experiments. We find that the BH accretion-related parameters are strongly dependent on the local BH mass function determination: higher normalizations and flatter high-mass slopes in the latter imply lower radiative efficiencies and mean Eddington ratios with a stronger redshift evolution. Additionally, the binary BH merger rate is estimated to be a fraction $\lesssim 10^{-1}$ of the galaxy merger rate derived from galaxy pairs counts by JWST, and constrained not to exceed the latter at $\gtrsim 2\sigma$. Relatedly, we highlight hints of a possible tension between current constraints on BH demographics and the interpretation of the nano-Hz GW background as predominantly caused by binary BH mergers. Specifically, we bound the latter’s contribution to $\lesssim 30-50\%$ at $\sim 3\sigma$, suggesting that either systematics in the datasets considered here have been underestimated so far, or that additional astrophysical/cosmological sources are needed to explain the residual part of the signal measured by PTA experiments.
@@ 1 Introduction
Relic black holes (BHs) with masses $M_{\bullet}\sim 10^{6}-10^{10}\,M_{\odot}$ are found ubiquitously at the center of massive galaxies in the local Universe (e.g., [1, 2, 3, 4]). These monsters are thought to have grown mainly by gaseous accretion onto a disk surrounding the BH (e.g., [5, 6]) that energizes the spectacular broadband emission of active galactic nuclei (AGNs) over a broad range of bolometric luminosities $L_{\rm AGN}\sim 10^{42}-10^{48}$ erg s-1. Such a paradigm has received a smoking gun confirmation by the Event Horizon Telescope [7, 8] via the imaging of the BH shadow caused by gravitational light bending and photon capture around the event horizon of M87 and Sgr A. Nevertheless, the evolutionary history of the supermassive BH population is still an open hot issue of modern astrophysics, as testified by the many studies conducted to address it in the last decades via numerical simulations (e.g., [9, 10, 11, 12, 13, 14, 15]), semi-analytic models (e.g., [16, 17, 18, 19, 20, 21]), and data-driven approaches (e.g., [22, 23, 24, 25, 26, 27, 28]).
@@
In fact, some uncertainties still hinder a robust reconstruction of the growth history of supermassive BHs, even on a population-averaged level. First of all, the initial conditions are poorly known. The simplest picture envisaging light seeds with $M_{\bullet}\lesssim 10^{2}\,M_{\odot}$ of stellar origin growing to the supermassive range by standard disk accretion has been challenged by the discovery of an increasing number of active BHs with masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$ already in place at appreciably high redshifts $z\gtrsim 7$, when the age of the universe was shorter than $0.8$ Gyr (e.g., [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]; for a recent review see [39]). This may require mechanisms able to rapidly produce heavy BH seeds $M_{\bullet}\sim 10^{3-5}\,M_{\odot}$, thus reducing the time needed to attain the final masses (see [40, 41, 42, 43, 44, 45, 46]).
@@
On the other end, the local demographics of the BH population still suffers from some systematics and possible biases. In fact, large samples of relic, quiescent BHs with precise mass determinations in the local Universe are not available. Thus the local BH mass function (i.e., the comoving number density of BHs as a function of mass) is usually estimated indirectly. A small sample with robust BH mass measurements is exploited to calibrate scaling relations of $M_{\bullet}$ with other host galaxy properties, like the stellar mass $M_{\star}$ or the stellar velocity dispersion $\sigma$ (e.g., [47, 48, 49, 50, 51, 52]). Such scaling relations are then employed to convert the statistics of a given galaxy property (e.g., the stellar mass function or the velocity dispersion function, determined on much larger samples of galaxies) into the BH mass function. However, this method is subject to systematics uncertainties that may drive the observed scaling relations away from the intrinsic ones [53, 54, 55, 56], hence potentially producing biased estimates of the local BH mass function.
@@
In addition, the intermediate stages of the BH evolution are also somewhat debated. For example, the accretion rates $\dot{M}_{\bullet}\propto L_{\rm AGN}/\epsilon$ and radiative efficiencies $\epsilon$ for the population of active BHs are loosely constrained, especially toward high-redshift [57, 58, 59, 60, 61, 62, 63]. The same apply to the distribution of Eddington ratio $\lambda\propto L_{\rm AGN}/M_{\bullet}$ (e.g., [64, 65, 66, 67, 37, 68]) or of its proxy $\lambda_{\star}\propto L_{X}/M_{\star}$ involving the X-ray luminosity $L_{X}$ and host stellar mass $M_{\star}$ (e.g., [69, 70, 71, 72]), which are only vaguely assessed for $z\gtrsim 4$. Even the redshift-dependent AGN luminosity function estimated via multi-wavelength surveys (e.g., [73, 74]) may suffer at $z\gtrsim 6$ from residual systematics on the adopted bolometric corrections and obscured fractions especially at the faint end (e.g., [75, 76, 77, 78, 79, 80, 81]). The overall picture has been made even more complex by the recent observations with JWST, which have identified at $z\gtrsim 4.5$ an abundant population of overmassive, X-ray under-luminous BHs (e.g., [82, 83, 84, 85, 86]); several of these are located in the so called ‘little red dots’ sources (see [87, 88, 89, 90, 91]), whose number density at $z\gtrsim 6$ may be up to two orders of magnitude higher than that of UV/optically selected AGNs [91, 92].
@@
Finally, the role of BH mergers, which constitute another relevant channel for the growth of supermassive BHs, is still under scrutiny. Binary BH merger events are thought to follow (maybe with some delay) the possible aggregation of the host galaxies, common in overdense environments especially toward low redshifts. However, so far such events have remained elusive and difficult to be corroborated, even indirectly, by observations of dual AGNs (e.g., [93, 94, 95]). Recently, the situation may have changed with the possible detection of the stochastic gravitational wave (GW) background in the nano-Hz frequency range by pulsar timing array (PTA) experiments [96, 97, 98, 99]. Although the statistical significance of the signal does not attain the $5\sigma$ threshold of confidence, and superimposed contributions from different physical mechanisms may prevent a clear-cut interpretation, such data offer a new observational handle on the BH growth process via binary mergers, as discussed in the literature with analytic approaches [100, 101, 102, 103, 104, 105, 106, 107, 108] and numerical simulations [109, 110, 111, 112].
In such a complex landscape, designing an ab-initio theoretical or numerical framework for the evolution of the supermassive BH population would be extremely challenging and perhaps not particularly illuminating. This is because the problem requires to self-consistently model many diverse processes cooperating on a huge range of temporal, spatial, and energy scales. This goes by far beyond our present physical understanding and computational capabilities, and in any case tends to produce outcomes strongly contingent on the chosen modeling, sub-grid recipes and spatial/temporal resolution (see [113, 114]). Therefore in this work we approach the issue via a complementary semi-empirical approach (see recent review by [115] and references therein), which is anchored by construction to the most reliable data available so far, takes into account observational/systematic uncertainties by transparently highlighting their impact on the results, and aims to minimize theoretical assumptions and associated parameterizations. In devising our data-driven methodology we have proceeded along the following consequential logical steps (formal details and further references will be provided in Section 2).
  • The first building block of our data-driven model is constituted by the observed bolometric AGN luminosity functions at different redshifts. This is a natural choice to start with, since AGN luminosity is the best tracer of gas accretion, which is expected to be the main growth channel of supermassive BHs. Moreover, the AGN luminosity function is a rather robust and well calibrated observable, especially at $z\lesssim 4$ where the aforementioned uncertainties are appreciably reduced; e.g., obscuration fraction and bolometric corrections are better established, while the little red dots population steadily decreases in number density to progressively become subdominant [116, 117]. Moreover, we exploit sound observational estimates of the Eddington ratios for active BHs at various redshift to design an educated parameterization of the related distribution. These ingredients are then employed as inputs into a continuity equation approach to derive the redshift-dependent demography of supermassive BHs as due to accretion.
  • Another crucial input parameter entering into the continuity equation is the radiative (plus kinetic) efficiency which regulates the amount of BH mass accumulated via gas accretion through cosmic time. In our framework, the efficiency and Eddington ratio distribution parameters are mostly constrained by the requirement to reproduce the local BH mass function. Thus the impact of systematics or biases in the latter is transparently highlighted by our model, in that different determinations of the local BH mass function will imply a corresponding variation of the efficiencies and Eddington ratios. Moreover, requiring the model to also reproduce large-scale BH clustering measurements will help to break any degeneracy among these quantities.
  • The outcome of the continuity equation is a unique solution for the amplitude, shape and evolution of the BH mass function as built by accretion. This in turn can be exploited as an input weight in a Smoluchowski coagulation equation approach to assess the role of binary supermassive BH mergers at population level. The main uncertainty in this approach is represented by the BH merger rate, which we anchor, in terms of a rescaling parameter, to the recent observational estimate of the redshift-dependent galaxy merger rate from galaxy pairs counts by JWST. In our framework, the overall impact of mergers is mainly constrained by the shape of the local BH mass function at high masses, and by the requirement to be consistent with (i.e. reproduce or at least not exceed) the nano-Hz GW background signal measured by PTA experiments.
  • We infer the parameters of our semi-empirical model by comparing it with the aforementioned datasets within a fully Bayesian framework. This approach enables a robust quantification of parameter degeneracies and statistical uncertainties. Consequently, we are able to rigorously assess the relative contributions of accretion and mergers to the evolutionary history of the supermassive BH population. Furthermore, we derive constraints on the fraction of the GW background that can plausibly be attributed to binary supermassive BH mergers. A key strength of our methodology is its ability to jointly constrain the evolution of the supermassive BH population by simultaneously fitting observational data on BH demographics, large-scale clustering, and the GW background.
In summary, in this work we employ a continuity plus Smoluchowski equation framework to unitarily describe the two main modes of growth for the supermassive BH population, namely gas accretion and binary mergers. We educatedly parameterize the main uncertainties associated to each BH growth channel, and infer joint constraints on these in a Bayesian setup from estimates of the local BH demographics, clustering, and measurements of the nano-Hz stochastic GW background. The plan of the paper is as follows: in Section 2 we describe the ingredients of our semi-empirical framework; in Section 3 we present our data analysis and discuss our results; in Section 4 we summarize our findings. Throughout this work we adopt the standard $\Lambda$CDM cosmology with parameters from the Planck mission [118].
2 Semi-empirical modeling

### v2

# Semi-empirical Framework of Supermassive Black Hole Evolution: Highlighting a possible tension between Demographics and Gravitational Wave Background

@@ Abstract
The evolution of the supermassive Black Hole (BH) population across cosmic times remains a central unresolved issue in modern astrophysics, due to the many noticeable uncertainties in the involved physical processes that span a huge range of spatial, temporal and energy scales. Here we tackle the problem via a semi-empirical approach with minimal assumptions and data-driven inputs. This is based on a continuity plus Smoluchowski equation framework that allows to unitarily describe the two primary modes of BH growth: gas accretion and binary mergers. Key quantities related to the latter processes are incorporated through educated parameterizations, and then constrained in a Bayesian setup from joint observational estimates of the local BH mass function, of the large-scale BH clustering, and of the nano-Hz stochastic gravitational wave (GW) background measured from Pulsar Timimg Array (PTA) experiments. We find that the BH accretion-related parameters are strongly dependent on the local BH mass function determination: higher normalizations and flatter high-mass slopes in the latter imply lower radiative efficiencies and mean Eddington ratios with a stronger redshift evolution. Additionally, the binary BH merger rate is estimated to be a fraction $\lesssim 10^{-1}$ of the galaxy merger rate derived from galaxy pairs counts by JWST, and constrained not to exceed the latter at $\gtrsim 2\sigma$. Relatedly, we highlight hints of a possible tension between current constraints on BH demographics and the interpretation of the nano-Hz GW background as predominantly caused by binary BH mergers. Specifically, we bound the latter’s contribution to $\lesssim 30-50\%$ at $\sim 3\sigma$, suggesting that either systematics in the datasets considered here have been underestimated so far, or that additional astrophysical/cosmological sources are needed to explain the residual part of the signal measured by PTA experiments.
@@ 1 Introduction
Relic black holes (BHs) with masses $M_{\bullet}\sim 10^{6}-10^{10}\,M_{\odot}$ are found ubiquitously at the center of massive galaxies in the local Universe (e.g., [Magorrian1998, Ferrarese2000, Gebhardt2000, Tremaine2002, Burke2025, Zou2025]). These monsters are thought to have grown mainly by gaseous accretion onto a disk surrounding the BH (e.g., [LyndenBell1969, Shakura1973]) that energizes the spectacular broadband emission of active galactic nuclei (AGNs) over a broad range of bolometric luminosities $L_{\rm AGN}\sim 10^{42}-10^{48}$ erg s-1. Such a paradigm has received a smoking gun confirmation by the Event Horizon Telescope [EHT2019, EHT2022] via the imaging of the BH shadow caused by gravitational light bending and photon capture around the event horizon of M87 and Sgr A. Nevertheless, the evolutionary history of the supermassive BH population is still an open hot issue of modern astrophysics, as testified by the many studies conducted to address it in the last decades via numerical simulations (e.g., [Hirschmann2014, RosasGuevara2016, Volonteri2016, Springel2018, Weinberger2018, Dave2019, Ni2022, Chakraborty2023, Zou2024]), semi-analytic models (e.g., [Kauffmann2000, Bower2006, Croton2006, Somerville2008, Dayal2019, Fontanot2020, Valiante2021, Ghodla2025]), and data-driven approaches (e.g., [Hopkins2006, Shankar2013, Aversa2015, Tucci2017, Carraro2020, Zhang2023, Bosi2025]).
@@
In fact, some uncertainties still hinder a robust reconstruction of the growth history of supermassive BHs, even on a population-averaged level. First of all, the initial conditions are poorly known. The simplest picture envisaging light seeds with $M_{\bullet}\lesssim 10^{2}\,M_{\odot}$ of stellar origin growing to the supermassive range by standard disk accretion has been challenged by the discovery of an increasing number of active BHs with masses $M_{\bullet}\gtrsim 10^{9}\,M_{\odot}$ already in place at appreciably high redshifts $z\gtrsim 7$, when the age of the universe was shorter than $0.8$ Gyr (e.g., [Mortlock2011, Wu2015, Venemans2018, Banados2018, Reed2019, Matsuoka2019, Yang2021, Wang2021, Farina2022, Dodorico2023]; for a recent review see [Fan2023]). This may require mechanisms able to rapidly produce heavy BH seeds $M_{\bullet}\sim 10^{3-5}\,M_{\odot}$, thus reducing the time needed to attain the final masses (see [Natarajan2014, Madau2014super, Mayer2019, Inayoshi2020, Boco2020, Volonteri2021, Pacucci2024]).
@@
At the other end, the local demographics of the BH population still suffers from some systematics and possible biases. In fact, large samples of relic, quiescent BHs with precise mass determinations in the local Universe are not available. Thus the local BH mass function (i.e., the comoving number density of BHs as a function of mass) is usually estimated indirectly. A small sample with robust BH mass measurements is exploited to calibrate scaling relations of $M_{\bullet}$ with other host galaxy properties, like the stellar mass $M_{\star}$ or the stellar velocity dispersion $\sigma$ (e.g., [Kormendy2013, McConnell2013, Reines2015, Davis2018, Sahu2019, Zhu2021]). Such scaling relations are then employed to convert the statistics of a given galaxy property (e.g., the stellar mass function or the velocity dispersion function, determined on much larger samples of galaxies) into the BH mass function. However, this method is subject to systematics uncertainties that may drive the observed scaling relations away from the intrinsic ones [Bernardi2007, Shankar2016, Shankar2019, Shankar2020], hence potentially producing biased estimates of the local BH mass function.
@@
In addition, the intermediate stages of the BH evolution are also somewhat debated. For example, the accretion rates $\dot{M}_{\bullet}\propto L_{\rm AGN}/\epsilon$ and radiative efficiencies $\epsilon$ for the population of active BHs are loosely constrained, especially toward high-redshift [Soltan1982, Salucci1999, Davis2011, Raimundo2012, Trakhtenbrot2017, Aird2022, Laloux2024]. The same applies to the distribution of Eddington ratio $\lambda\propto L_{\rm AGN}/M_{\bullet}$ (e.g., [Kollmeier2006, Vestergaard2009, Willott2010, Kelly2013, Farina2022, Wu2022]) or of its proxy $\lambda_{\star}\propto L_{X}/M_{\star}$ involving the X-ray luminosity $L_{X}$ and host stellar mass $M_{\star}$ (e.g., [Bongiorno2016, Georgakakis2017, Aird2018, Carraro2022]), which are only vaguely assessed for $z\gtrsim 4$. Even the redshift-dependent AGN luminosity function estimated via multi-wavelength surveys (e.g., [Hopkins2007LF, Shen2020]) may suffer at $z\gtrsim 6$ from residual systematics on the adopted bolometric corrections and obscured fractions especially at the faint end (e.g., [Lusso2012, Duras2020, Ueda2014, Buchner2015, Ananna2019, Laloux2023, Andonie2024]). The overall picture has been made even more complex by the recent observations with JWST, which have identified at $z\gtrsim 4.5$ an abundant population of overmassive, X-ray under-luminous BHs (e.g., [Ubler2023, Furtak2023, Harikane2023, Kokorev2023, Stone2024]); several of these are located in the so called ‘little red dots’ sources (see [Kocevski2023, Matthee2024, He2024, Lai2024, Kokorev2024]), whose number density at $z\gtrsim 6$ may be up to two orders of magnitude higher than that of UV/optically selected AGNs [Kokorev2024, Akins2025].
@@
Finally, the role of BH mergers, which constitute another relevant channel for the growth of supermassive BHs, is still under scrutiny. Binary BH merger events are thought to follow (maybe with some delay) the possible aggregation of the host galaxies, common in overdense environments especially toward low redshifts. However, so far such events have remained elusive and difficult to be corroborated, even indirectly, by observations of dual AGNs (e.g., [Comerford2013, Silverman2020, Pfeifle2025]). Recently, the situation may have changed with the possible detection of the stochastic gravitational wave (GW) background in the nano-Hz frequency range by pulsar timing array (PTA) experiments [Agazie2023, EPTA2024, Reardon2023, Xu2023]. Although the statistical significance of the signal does not attain the $5\sigma$ threshold of confidence, and superimposed contributions from different physical mechanisms may prevent a clear-cut interpretation, such data offer a new observational handle on the BH growth process via binary mergers, as discussed in the literature with analytic approaches [Wyithe2003, Sesana2008, Kulier2015, Sesana2016, Dascoli2018, Chen2019, Simon2023, SatoPolito2024, SatoPolito2025] and numerical simulations [Salcido2016, Kelley2017, Siwek2020, Chen2025].
To address the numerous facets and issues of such a complex landscape, three main complementary approaches are possible and have been largely exploited in the last decades: hydrodynamical simulations (for a review, see [Naab2017]), semi-analytic models (for a review, see [Somerville2015rev]), and semi-empirical frameworks (for a review, see [Lapi2025]). Hydrodynamical simulations allow to address the galaxy and BH evolution process in fine detail, as they can treat simultaneously dark matter, gas and stars. However, despite the recent increase in resolution and speed (mainly thanks to emulators and ’genetic’ codes), many of the relevant physical processes still constitute sub-grid physics, while a detailed exploration of the parameter space is often limited by rather long computational times. Semi-analytic models are based on dark matter merger trees extracted from NN-body simulations (or generated via Monte Carlo procedures gauged on them), while the physics inside dark halos is modeled via parametric expressions set on (mainly) local observables. These models are less computationally expensive than hydrodynamical simulations and enable to test different prescriptions more easily, though the considerable number of fudge parameters describing the underlying physics can lead to degenerate solutions and restrict somewhat the predictive power, especially toward high redshift (e.g., [Seeyave2025]). Concerning the evolution of the (super)massive BH population, ab-initio models struggle somewhat in reproducing coherently the wealth of observational data and in elucidating the different role of the main physical mechanisms at work. This is mainly because the problem requires to self-consistently model many diverse processes cooperating on a huge range of temporal, spatial, and energy scales. Such a task goes by far beyond our present physical understanding and computational capabilities, and in any case tends to produce outcomes strongly contingent on the chosen modeling, sub-grid recipes and spatial/temporal resolution (see [Donnari2021, Habouzit2022, Crain2023]).
The situation calls for a reverse-engineering approach that inverts the traditional methodology: rather than testing ab-initio model predictions against observations, one can construct semi-empirical frameworks that extract physical insights directly from data with minimal theoretical assumptions. While these frameworks may not capture the full theoretical richness of ab-initio models or provide complete physical narratives, their strength lies in delivering robust, model-independent constraints on key observables and providing clear answers to specific, well-posed questions. Moreover, by empirically linking different observables, they can check for possible inconsistencies or tensions among distinct datasets. Such data-driven models serve not to replace more sophisticated semi-analytic and numerical approaches, but are meant as complementary tools to provide independent constraints that can break existing degeneracies and guide theoretical developments toward more physically-motivated parameterizations.
[…]
2 Methodology
In this Section we first provide a qualitative overview of our data-driven methodology, which is also schematically illustrated in Figure 1, while formal details and references can be found in the following subsections.
  • The core tool of the method is constituted by the continuity equation. This provides an effective way of deriving the statistical evolution of the (super)massive BH population from data-driven inputs. Specifically, the continuity equation is a partial differential equation which evolves timestep by timestep the BH mass function across cosmic time, taking into account the averaged growth of the BH population due to gas accretion and mergers.
  • The component related to gas accretion is described by a drift term in the continuity equation. It requires two inputs: the observed redshift-dependent bolometric AGN luminosity function $N(\log L_{\rm AGN},t)$, which provides the statistics of active BHs emitting at a given luminosity $L_{\rm AGN}$; a parameterized form of the Eddington ratio distribution $P(\log\lambda)$, which basically informs on how BHs of different masses are contributing to the different luminosity bins. The parameters involved are three: two ($\log\lambda_{0}$ and $\xi$) describe the local normalization and redshift evolution for the mean of the Eddington ratio distribution; the other one is the radiative efficiency $\epsilon$, which allows to convert AGN luminosity into BH accretion rates, eventually needed to compute BH masses.
  • The component related to mergers is described by a Smoluchowski coagulation (kind of diffusion) term in the continuity equation. It requires one input and two additional parameters: in the reasonable assumption that BH mergers follow galaxy mergers, one parameter $f_{\bullet\bullet\rightarrow\bullet}$ is a rescaling factor for the the redshift-dependent normalization of the BH merger rate with respect to the galaxy merger rate as estimated from galaxy pair counts, which constitutes the observational input (but we will explore variants where the BH merger rate is maximized by adopting $f_{\bullet\bullet\rightarrow\bullet}\approx 1$); another parameter $\eta$ specifies a possible additional dependence of the BH binary merger rate on the BH mass.
  • The direct output of the continuity equation is a unique solution for the amplitude, shape and evolution of the BH mass function $N(\log M_{\bullet},t)$ as built by accretion and mergers. The BH mass function at $z\approx 0$ can be directly compared with independent observational determinations from local galaxy statistics and BH-to-galaxy scaling relations.
[…]
  • Another derived outcome is the stochastic GW background from BH mergers in the nano-Hz frequency range. This can be straightforwardly computed from the outputs of the continuity equation and in particular from the redshift-dependent BH mass function and binary BH merger rate (related to the Smoluchowski coagulation term). The resulting GW background amplitude can then be directly compared with the current observational estimates from PTA experiments.
  • The aforementioned parameters of our semi-empirical framework (namely, the radiative efficiency of BH accretion, the mean of the Eddington ratio distribution and its redshift evolution, the normalization and mass dependence of the BH merger rate, and the scatter in the BH vs. halo mass relation) are inferred by comparing the direct output and the derived outcomes with independently measured datasets within a fully Bayesian setup. This approach enables a robust quantification of parameter degeneracies and statistical uncertainties. A key strength of our framework is its ability to jointly constrain the evolution of the supermassive BH population by simultaneously fitting observational data on BH demographics, large-scale clustering, and the GW background. We will discuss below how such a combination of independent observables enables us to highlight a possible tension between the amount of GW background due to binary BH merger and the measured signal from PTA experiments, which could instead be missed by BH merger models that attempt to only fit the GW background ignoring all the other demographic indicators.
In summary, the continuity equation is a flexible framework that by integrating a few observational inputs (e.g., AGN luminosity functions) yields a clear direct output (e.g., BH mass function) and some derived outcomes (e.g., BH clustering, PTA signal) via minimal physical assumptions (e.g., BHs grow by gas accretion and binary mergers) and a limited number of parameters (six in total, see above). The direct output and derived outcomes are then compared with independent datasets to obtain a best estimate of the parameters, which ultimately yields insights into the physical evolution of the BH population. Such a data-driven approach bypasses the complex and uncertain modeling of the baryonic physics associated to galaxies within dark matter halos via semi-analytic recipes or hydrodynamical simulations. Moreover, since the number of observables and parameters are comparable, it is not prone to large degeneracies but only affected by possible systematics. All in all, our semi-empirical framework can robustly assess the relative contributions of accretion and mergers to the evolutionary history of the supermassive BH population, whilst constraining the fraction of the GW background that can plausibly be attributed to supermassive BH mergers.
2.1 The continuity equation